OpenStudy (anonymous):

y= -3x^2 + 5x - 7 to vertex form.

OpenStudy (anonymous):

first coordinate of vertex is \[-\frac{b}{2a}=-\frac{5}{2\times (-3)}=\frac{5}{6}\]\]

OpenStudy (anonymous):

second coordinate is whatever you get when \[x=\frac{5}{6}\] i get \[-\frac{59}{12}\]

OpenStudy (anonymous):

so "vertex form" is \[y=-3(x-\frac{5}{6})^2-\frac{59}{12}\]

OpenStudy (anonymous):

thnx